There are 2856 three's in the product of 51 and a number. What is the number?
Answer by Theo(13342)

your equation is:
51 * x = 3 * 2856
solve for x to get:
x = (3 * 2856) / 51
you get:
x = 168


to confirm do the following:
168 * 51 = 8568
divide that by 3 and you get 2856.